If one vehicle is going eastbound at 20 mph, and another vehicle is traveling westbound at 55 mph, how many seconds would it take the car going eastbound to turn around and catch up to the westbound car ?
Answer by khwang(438)

Missing conditions:
Unless there is some information to have the difference of their apart at
some time, we cannot solve this problem.
Usually,the time of catching up = D(distance) / (V2 - V1)
= D/(55 -20) = D/35.
Also, this is a pretty easy question. I hope you can solve it by
yourself without posting again after read this hint and get the enough
information about the distance.